OpenStudy (anonymous):

If f(x)= 2x 2 + 3, then f(x - 1)= 2x 2 + 4x + 5. how is this true

OpenStudy (kainui):

Everywhere you see "x" in f(x) you replace with "x-1" to find f(x-1).\[f(x)=2x^2+3\]\[f(x-1)=2(x-1)^2+3\] see? Now you just multiply the polynomial out.

OpenStudy (kainui):

Are you sure you typed the question correctly?

OpenStudy (anonymous):

it was copied correctly